OpenStudy (daniel.ohearn1):

How could one derive the following inequality from the Riemann Hypothesis?

OpenStudy (daniel.ohearn1):

pi(x)+li(x)-li(2) < sqrt(x)ln(x)

OpenStudy (daniel.ohearn1):

@Zarkon @Zenmo @zepdrix @Cublade @VladthePro @myininaya @FaireGaming @Gonin12 @heyyyyitsholly @jasminewhitaker1111 @kitty10280 @LunyMoony @Qwertty123 @wolf1728 @radar @TheSmartOne @ilovemath14 @PokieOkie

OpenStudy (ilovemath14):

idk this type of math im sorry

OpenStudy (daniel.ohearn1):

li(x)= \[\int\limits_{0}^{x}\frac{ 1 }{ \log(t) }dt \]

OpenStudy (daniel.ohearn1):

Then we have our friend the Zeta function: \[\zeta(s)=\prod\frac{ 1 }{ 1-p ^{-s}}\]